OpenStudy (anonymous):

Suppose the graph of y=f(x) contains the point (9,7). Then the point (1,7) must be on the graph of which of the following?

OpenStudy (anonymous):

(a) y=9f(x) (b) y=f(9x) (c) y=f(x-8) (d) y=f(1/9x) (e) y=f(x)-8

OpenStudy (anonymous):

i guess b

OpenStudy (anonymous):

since \[f(9)=7\] and therefore \[f(9\times 1)=f(9)=7\]

OpenStudy (anonymous):

B is right f(1) = 7 f(1x9) = 7